# Large Intelligent Surfaces with Low-End Receivers: From Scaling to Antenna and Panel Selection

### v2

@@ Abstract
We analyze the performance of large intelligent surface (LIS) with hardware distortion at its RX-chains. In particular, we consider the memory-less polynomial model for non-ideal hardware and derive analytical expressions for the signal to noise plus distortion ratio after applying maximum ratio combining (MRC) at the LIS. We also study the effect of back-off and automatic gain control on the RX-chains. The derived expressions enable us to evaluate the scalability of LIS when hardware impairments are present. We also study the cost of assuming ideal hardware by analyzing the minimum scaling required to achieve the same performance with a non-ideal hardware. Then, we exploit the analytical expressions to propose optimized antenna selection schemes for LIS and we show that such schemes can improve the performance significantly. In particular, the antenna selection schemes allow the LIS to have lower number of non-ideal RX-chains for signal reception while maintaining a good performance. We also consider a more practical case where the LIS is deployed as a grid of multi-antenna panels, and we propose panel selection schemes to optimize the complexity-performance trade-offs and improve the system overall efficiency.
Index Terms:
Hardware Distortion, Large Intelligent Surface, Massive MIMO, Panel Selection, Receive Antenna Selection.
@@ I Introduction
Large Intelligent Surfaces (LISs) have emerged as one of the next major steps in the future development of multiple-input multiple-output (MIMO) communication systems [1, 2, 3]. While LISs are sometimes considered to be a scaled up version of the widely known massive MIMO systems, they have some unique properties which distinguishes them from the massive MIMO systems. For instance, their physical size can be compared to the user-equipment (UE) distance to the LIS, which introduces near-field effects in the wireless channel. As a result, the common assumptions about the massive MIMO channels are no longer valid and new effects emerge. This requires different channel models, and more investigation into the already known facts about the transmit and receive schemes in massive MIMO systems [4].
[…]
@@
The deployment of LISs is envisioned to be much more challenging because of the enormous leap in the number of transceiver chains, and the associated processing complexities. One important challenge is the cost efficiency of the whole system, where hundreds to thousands of transceiver chains are to be deployed, which may force systems designers to consider the use of inexpensive hardware components in each of these transceiver chains [7]. Therefore, it is of high importance to study and mitigate the effects of hardware distortion in transceiver chains when scaling up LIS to allow the use of cost-efficient transceiver chains. There has been previous works on the design and optimizing IRS and RIS based system with hardware impairment [8, 9, 10]. However, these works have no direct applicability to our study due to the intrinsically different system model associated to passive reflective surfaces compared to large transceiver technologies as LIS.
The vision for future antenna arrays in the next generation of wireless communication systems highly depends on deploying arbitrarily large arrays due to the compelling gains in terms of multiplexing and beamforming performance through increasing spatial resolution [11]. With traditional design approaches, implementing such a huge array may not be economically favored for operators and vendors, since each antenna is typically equipped with an analog and a digital front-end (AFE and DFE) [6]. The requirements for each transceiver chain may lead to unreasonable implementation costs [12]. On the other hand, the power consumption of each transceiver can also become a bottleneck, especially if there are tight requirements on the non-linearity of amplifiers and the out-of-band emissions. AFEs are considered to be one of the main sources of hardware distortion in receivers [13, 14, 15]. Not only the non-linearity effect of AFEs are important to consider, but also the limitations in the total power consumption of the receiver chain is of great importance. Linearity requirements in the RX-chains can result in a huge leap in power consumption of the whole system [16], which is not favorable neither in terms of cost nor energy efficiency.
@@
To deal with these challenges, we need to make the receiver design as efficient as possible, from a system perspective. In general, the path to follow is to optimize the signal processing schemes and system designs while maintaining the hardware quality at a minimum level. In other words, we want to get the most gain from each transceiver chain while limiting the implementation costs, e.g., by using inexpensive hardware components. In [17] and [18], we have proposed approaches to address such challenges in massive MIMO systems, mainly by optimizing the per-antenna digital pre-distortion (DPD) resources. Another approach is to perform antenna selection methods and only use a portion of the array for transmission and reception. By doing so, we only use antennas which have a significant contribution to the performance and thereby increase energy efficiency.
@@
In this paper, we focus on optimizing the performance of LIS with non-ideal RX-chains. We first analyze the signal to distortion-plus-noise ratio (SNDR) with the purpose of studying the scaling behavior and asymptotic limits of LIS with hardware distortion. Then, we propose receive antenna selection schemes for a LIS with non-ideal AFEs. Optimization problems are defined and solved to illustrate the importance of performing antenna selection when scaling up LIS. We show that selecting antennas with the strongest channels is not always the optimal solution. We then focus on more practical cases where the LIS is implemented as a grid of panels, and transform the antenna selection problem into a panel selection problem. Low-complexity closed-form sub-optimal solutions are proposed for the panel selection. We also show that, by adopting such antenna selection and panel selection schemes, we can improve the system performance significantly for a fixed receive-chain hardware quality.
@@ II System Model
In, [4], the effect of considering near-field properties in the channel model is studied, including the variation of distance from each antenna element to the UE, different effective area, and different polarization losses due to different angle of arrivals. The results in [4] imply that only considering the first two effects or considering all the three effects , even if the number of antenna elements grows extremely large, have negligible difference. Therefore, we only consider the first two effects and adopt the near-field LOS channel model as considered in [1]. The narrow-band channel $h(x,y)=|h(x,y)|~{}\text{exp}\left(-j\phi(x,y)\right)$ between the UE and a point on the LIS with coordinates $(x,y,0)$ has amplitude given by
@@
where $\lambda$ is the wavelength. For the $n$’th antenna element, located at $(x_{n},y_{n},0)$ with effective area $A$ small enough such that $|h(x,y)|^{2}$ is almost fixed inside its effective area [^1] This assumption is valid in most practical MIMO scenarios since the effective area is in the orders of $\frac{\lambda^{2}}{4}$ and the UE is in the far-field of each individual antenna element, i.e. $d\gg\frac{\lambda}{2}$., the channel gain is $|h_{n}|^{2}=A|h(x_{n},y_{n})|^{2}$.
While this form of deployment is a common vision for the LIS implementation in the future of wireless networks [20], it is not favorable with current technology from a practical point of view [21]. A potentially more favorable option, without losing practicality and cost efficiency is to implement the LIS as grid of panels, each equipped with some equally distanced antenna elements [6]. Fig. 2 illustrates an example of panel configuration of a LIS on the $x$-$y$ plane.
@@ III SNDR Characterization
where $\bm{G}=\text{diag}\left(g_{1},\dots,g_{N}\right)$, $\bm{C_{\eta\eta}}=\text{diag}\left(C_{1},\dots,C_{N}\right)$, with $g_{n}$ and $C_{n}$ corresponding to the Busssgang gain and distortion power for the $n$’th antenna as given in (7) and (8), respectively. For analytical tractability and based on the results from [27], we have assumed that the distortion correlation among the antennas are negligible, i.e. $\bm{C_{\eta\eta}}$ is diagonal.
The LIS applies a combing vector $\bm{v}$ to the received signal $\bm{r}$ to equalize the transmitted signal. Maximum Ratio Combining (MRC) is an attractive option in LIS scenarios because of its simplicity and reasonable performance, since it still allows us to exploit the available spatial degrees of freedom[1]. In our scenario, the MRC vector may be expressed as $\bm{v}^{T}=\tilde{\bm{h}}^{H}/\|\tilde{\bm{h}}\|$ where $\tilde{\bm{h}}=\bm{G}\bm{h}$, can be interpreted as the effective channel, which includes the physical channel and the multiplicative hardware distortion effect. In fact, the LIS would only be able to estimate the effective channel $\tilde{\bm{h}}$ from the uplink UE pilots since the signals received during channel estimation are also affected by the hardware distortion. We assume that the LIS has a perfect estimate of $\tilde{\bm{h}}$ so that we can isolate the effect of hardware distortion from that of channel estimation imperfections. Note that the estimation error would only correspond to an uncorrelated additive term, which would effectively not affect our analysis on hardware distortion.
@@
The signal, distortion, and noise power can be calculated from (16) considering that the distortion term is uncorrelated with the signal and has covariance $\bm{C_{\eta\eta}}$. The SNDR is therefore given by
@@
[…]
@@ IV LIS Antenna Selection
The contribution of each antenna to the signal and distortion power depends on the received power to that individual antenna. Antennas receiving more power may contribute more to the useful signal power at the cost of introducing more distortion. In Section III, we have seen how the receiver array size, given by the LIS radius, affects the performance of the whole system. We assumed that the area for signal reception was selected as a disk around the center of the array, assuming that UE was also aligned with the central perpendicular line. The question that arises here would be: Can we improve the signal reception performance, i.e. SNDR, by selecting the antennas from another region? In other words, can we get an improvement from considering a reception area with a different shape than a disk with a center aligned with the UE position? This question can be formulated as an antenna selection problem, where we consider a LIS with $N\gg 1$ elements, and a resource constraint which forces the LIS to only use $N_{\text{max}}$ of the antenna elements for signal reception.
@@
$\max_{z_{1},\dots,z_{N}}~{}$ $\frac{P\sum_{n=1}^{N}z_{n}\left|\tilde{h}_{n}\right|^{2}}{\frac{\sum_{n=1}^{N}z_{n}C_{n}\left|\tilde{h}_{n}\right|^{2}}{\sum_{n=1}^{N}\left|\tilde{h}_{n}\right|^{2}}+\sigma^{2}},$ (41)
@@
Let us define the set of selected antennas as =$\mathcal{F}=\{i\in\mathbb{N}~{}|~{}r<x_{i}^{2}+y_{i}^{2}<R\}$. Therefore, we have $z_{i}=1$ if $i\in\mathcal{F}$. By leveraging Lemma III.1, the following approximation then holds for the numerator in (41),
@@
$\frac{\sum_{n=1}^{N}z_{n}C_{n}\left|\tilde{h}_{n}\right|^{2}}{\sum_{n=1}^{N}\left|\tilde{h}_{n}\right|^{2}}=\frac{P\kappa\sum_{n\in\mathcal{F}}\left|{h}_{n}\right|^{4}}{\sum_{n\in\mathcal{F}}|h_{n}|^{2}}$ (43)
@@
which has the benefit of being a continuous function of only two variables, and it has a reduced complexity over (41) when solved by numerical methods. An algorithm to solve this problem numerically is introduced in Appendix B.
@@ V LIS Panel Selection
$\max_{z_{n}}~{}$ $\frac{PM\sum_{n=1}^{N_{p}}z_{n}\left|\tilde{h}_{n}\right|^{2}}{\frac{\sum_{n=1}^{N_{p}}z_{n}C_{n}\left|\tilde{h}_{n}\right|^{2}}{\sum_{n=1}^{N_{p}}\left|\tilde{h}_{n}\right|^{2}}+\sigma^{2}},$ (46)
@@
$0\displaystyle 2\beta|{a}_{3}|^{2}\rho^{4}+4\alpha|{a}_{3}|^{2}\rho^{3}-2\beta\sigma^{2}\rho-\alpha\sigma^{2}=$ 0 . (50)
@@
[…]

### v3

@@ Abstract
Feasibility of the promising large intelligent surface (LIS) concept, as well as its scalability, relies on the use of low-cost hardware components, raising concerns about the effects of hardware distortion. We analyze LIS systems with receive-chain (RX-chain) hardware distortion, showing how it may limit performance gains when scaling up these systems. In particular, using the memory-less polynomial model, analytical expressions are derived for the signal to noise plus distortion ratio (SNDR) after applying maximum ratio combining (MRC). We also study the effect of back-off and automatic gain control on the RX-chains. The derived expressions enable us to evaluate the scalability of LIS when hardware impairments are present. The cost of assuming ideal hardware is further analyzed by quantifying the minimum scaling required to achieve the same performance with non-ideal hardware. The analytical expressions derived in this work are also used to propose practical antenna selection schemes for LIS, and we show that such schemes can improve the performance significantly leading to increased energy efficiency. Specifically, by turning off RX-chains with lower contribution to the post-MRC SNDR, we can reduce the energy consumption while maintaining performance. We also consider a more practical scenario where the LIS is deployed as a grid of multi-antenna panels, and we propose panel selection schemes to optimize the complexity-performance trade-offs and improve the system overall efficiency.
@@ I Introduction
Large Intelligent Surfaces (LISs) have emerged as one of the next major steps in the future development of multiple-input multiple-output (MIMO) communication systems [1, 2, 3]. While large intelligent surfaces are sometimes considered to be a scaled up version of the widely known massive MIMO systems, they have some unique properties which distinguishes them from the massive MIMO systems. For instance, their physical size can be compared to the user-equipment (UE) distance to the LIS, which introduces near-field effects in the wireless channel. As a result, the common assumptions about the massive MIMO channels are no longer valid and new effects emerge. This requires different channel models, and more investigation into the already known facts about the transmit and receive schemes in massive MIMO systems [4]. The superiority of LIS in exploiting the available spatial degrees of freedom in the environment has been established in [1], for the line-of-sight (LOS) scenario, and further confirmed by the results in [5, 6] for more general propagation environments.
The LIS technology was initially introduced in [1] as a large continuous surface of electromagnetically active material in a LOS channel model. While having LIS as a continuous surface is useful to conceptually understand the system, it does not seem realistically implementable with current technology in the near future. Instead, a dense antenna array deployed throughout a large surface is more realistic, which can also be interpreted as a sampled version of the continuous LIS model. It has been shown that, if sampling is dense enough, specifically half-wavelength spatial sampling, the continuous and sampled versions of LIS have the same performance in terms of available spatial degrees of freedom due to the spatially-band-limited nature of the channels [1, 6]. More practical implementations of LIS consider dividing the LIS surface into dense multi-antenna panels with a lower number of antennas, which is a more flexible and scalable design with reduced system complexity [7]. What we refer to as LIS in this manuscript, corresponds to a base station (BS) technology consisting of an active sampled surface, which should not be confused with a reconfigurable intelligent surface (RIS) or an intelligent reflecting surface (IRS).
@@
The deployment of LISs is envisioned to be much more challenging than conventional BS technology because of the enormous leap in the number of transceiver chains, and the associated processing complexities. While this is motivated by the compelling gains in terms of multiplexing and beamforming performance through increasing spatial resolution [8], the cost-efficiency of the whole system can be a great challenge due to unreasonable implementation overheads associated with deploying hundreds to thousands of transceiver chains [9]. This may force system designers to consider the use of inexpensive hardware components in each of these transceiver chains [10], which can introduce distortion at the output of the transceivers. Therefore, it is of high importance to study and mitigate the effects of hardware distortion in transceiver chains when scaling up LIS to allow the use of cost-efficient transceiver chains.
There has been previous works on the design and optimization of RIS/IRS-based systems with hardware distortion [11, 12, 13, 14]. For example, [12] investigates channel estimation in IRS-assisted MISO systems under imperfect hardware, which is modeled by additive distortion. [13] analyzes the achievable rate of IRS-assisted wireless systems considering hardware impairments at both the transceivers and the reflecting surface, also modeled by additive distortion. It further proposes phase shift optimization strategies that mitigate these impairments. A complete survey on these approaches can be found in [11]. However, the literature on RIS/IRS with hardware imperfections has no direct applicability to our study due to the intrinsically different system model associated to passive reflective surfaces compared to active transceiver technologies as LIS, which is the main focus of this work.
@@
Transceiver AFEs non-linearity is considered to be one of the main sources of hardware distortion in receivers [15, 16, 17]. It has been shown that deploying AFEs with less non-linearity may lead to major leaps in power consumption [18]. Therefore, to deal with the challenge of implementing LIS in a cost- and energy-efficient manner, it is of great importance to adopt specific schemes to compensate the effect of employing highly non-linear AFEs. A potential path to follow is to optimize the signal processing schemes and system design while maintaining the hardware quality at a minimum level. In other words, we want to get the most gain from each transceiver chain while limiting the implementation costs, e.g., by using inexpensive hardware components with limited power consumption. In [19] and [20], approaches were proposed to address such challenges in massive MIMO systems, mainly by optimizing the per-antenna digital pre-distortion (DPD) resources. Another approach is to implement antenna selection schemes and only use a portion of the array for transmission and reception. By doing so, we only use antennas which have a significant contribution to the performance and thereby increase energy efficiency.
@@
In this paper, we focus on optimizing the performance of LIS with non-ideal RX-chains. We first analyze the signal to noise plus distortion ratio (SNDR) with the purpose of studying the scaling behavior and asymptotic limits of LIS with hardware distortion. Then, we propose receive antenna selection schemes for a LIS with non-ideal AFEs. Optimization problems are defined and solved to illustrate the importance of performing antenna selection when scaling up LIS. We show that selecting antennas with the strongest channels is not always the optimal solution. We then focus on more practical cases where the LIS is implemented as a grid of panels, and transform the antenna selection problem into a panel selection problem. Low-complexity closed-form sub-optimal solutions are proposed for the panel selection. We also show that, by adopting such antenna selection and panel selection schemes, we can improve the system performance significantly for a fixed receive-chain hardware quality.
@@ II System Model
In, [4], the effect of considering near-field properties in the channel model is studied, including the variation of distance from each antenna element to the UE, different effective area, and different polarization losses due to different angle of arrivals. The results in [4] imply that only considering the first two effects or considering all the three effects , even if the number of antenna elements grows extremely large, have negligible difference. Therefore, we only consider the first two effects and adopt the near-field LOS channel model as considered in [1]. Note that in LIS systems, the probability of having a strong LOS component is higher than in conventional massive MIMO due to the significantly larger aperture covered by these systems [1]. This effect is even further pronounced at higher frequencies, where reflected waves suffer greater attenuation [22, 23]. The amplitude of the narrow-band channel $h(x,y)=|h(x,y)|~\text{exp}\left(-j\phi(x,y)\right)$ between the UE and a point on the LIS with coordinates $(x,y,0)$ may then be expressed as
@@
where $\lambda$ is the wavelength. For the $n$’th antenna element, located at $(x_{n},y_{n},0)$ with effective area $A$ small enough such that $|h(x,y)|^{2}$ is almost constant throughout the area, [^1] This assumption is valid in most practical MIMO scenarios since the effective area is in the orders of $\frac{\lambda^{2}}{4}$ and the UE is in the far-field of each individual antenna element, i.e. $d\gg\frac{\lambda}{2}$. the channel gain is $|h_{n}|^{2}=A|h(x_{n},y_{n})|^{2}$.
While deploying LIS as a co-located array of antennas is a common vision for future of wireless networks [24], it may not be favorable in terms of flexibility and scalability [25]. A potentially more favorable option, without losing practicality and cost efficiency is to implement the LIS as grid of panels, each equipped with a smaller number of equispaced antenna elements [7]. Fig. 2 illustrates an example of panel configuration of a LIS on the $x$-$y$ plane.
@@
Remark: This work focuses on a single-UE scenario so as to isolate the impact of hardware distortion in LIS systems, facilitating a clearer understanding of this particular effect. However, the main results can be extended to multi-UE scenarios, specially in hardware-distortion-limited cases where they may be directly applied. For example, in the case of multi-user scenarios with sufficiently spaced users, each of the UEs may be associated to non-overlapping active LIS areas around the respective central perpendicular lines, which is still reasonable given the large dimensionality of LIS systems.
@@ III SNDR Characterization
where $\bm{G}=\text{diag}\left(g_{1},\dots,g_{N}\right)$, with $g_{n}$ corresponding to the Busssgang gain for the $n$’th antenna as given in (7), and $\bm{\eta}$ is the vector containing additive distortion at the LIS, with covariance matrix $\bm{C_{\eta\eta}}$. In general, $\bm{C_{\eta\eta}}$ is a non-diagonal matrix. However, the distortion correlation effect among antennas can be neglected in some cases, resulting in over-estimating the SNDR, which is often considered in the literature [31]. The effect of disregarding the distortion correlation is studied in Section VI-C.
The LIS applies a combing vector $\bm{v}$ to the received signal $\bm{r}$ to equalize the transmitted signal. Maximum Ratio Combining (MRC) is an attractive option in LIS scenarios because of its simplicity and reasonable performance, since it still allows us to exploit the available spatial degrees of freedom[1]. In the single-UE case without hardware distortion, MRC further provides the optimal equalization approach in terms of capacity, since it maximizes the post-processed SNR. In our scenario, the MRC vector may be expressed as $\bm{v}^{T}=\tilde{\bm{h}}^{H}/\|\tilde{\bm{h}}\|$ where $\tilde{\bm{h}}=\bm{G}\bm{h}$, can be interpreted as the effective channel, which includes the physical channel and the multiplicative hardware distortion effect. In fact, the LIS would only be able to estimate the effective channel $\tilde{\bm{h}}$ from the uplink UE pilots since the signals received during channel estimation are also affected by the hardware distortion. We assume that the LIS has a perfect estimate of $\tilde{\bm{h}}$ so that we can isolate the effect of hardware distortion from that of channel estimation imperfections. Note that the estimation error would only correspond to an uncorrelated additive term, which would effectively not affect our analysis on hardware distortion.
@@
The signal, distortion, and noise power can be calculated from (16) considering that the distortion term is uncorrelated with the signal and has covariance $\bm{C_{\eta\eta}}$. The instantaneous SNDR is therefore given by
$\gamma=\frac{P\tilde{\bm{h}}^{H}\tilde{\bm{h}}}{\frac{\tilde{\bm{h}}^{H}\bm{C_{\eta\eta}}\tilde{\bm{h}}}{\tilde{\bm{h}}^{H}\tilde{\bm{h}}}+\sigma^{2}}.$ (17)
If we assume that the distortion correlation among the antennas are negligible, i.e., $\bm{C_{\eta\eta}}$ is diagonal, the instantaneous SNDR can be simplified, which results in more analytical tractability for studying the impact of hardware distortion on LIS systems. We analyze the effect of this assumption in Section VI-C and show that it does not invalidate the conclusions drawn regarding scaling up of LIS systems with hardware distortion. Moreover, it leads to significant simplification of the proposed methods, at the cost of only a minor performance loss. Assuming diagonal $\bm{C_{\eta\eta}}$, i.e., $\bm{C_{\eta\eta}}=\text{diag}\left(C_{1},\dots,C_{N}\right)$, where $C_{n}$ corresponds to the distortion power for the $n$’th antenna, the instantaneous SNDR is simplified to
@@
As remarked before, most of the results and conclusions of this paper can be extended to a multi-user scenario, especially if the UEs are well-separated, or when the hardware distortion is the main source of error compared to the inter-user interference. In a multi-user scenario, the SNDR in (18) and (21) would contain an additional term in the denominator corresponding to the inter-user interference.
@@
$\sum_{n=1}^{N}|h_{n}|^{2}$ $\approx\iint\limits_{\mathcal{S}}|h(x,y)|^{2}dxdy$
$=\int_{\phi=0}^{2\pi}\int_{r=0}^{R}|h(r)|^{2}rdrd\phi,$ (22)
@@ IV LIS Antenna Selection
As we have seen in previous sections, and verified through numerical results in Section VI, hardware distortion can be a major limiting factor when scaling up LIS systems. Introducing more back-off, employing higher quality hardware components, or using more transceiver chains can be seen as alternatives to improve system performance affected by hardware distortion [18, 34]. However, all these solutions result in lower energy efficiency (EE) and cost-efficiency of the LIS systems, which can pose a burden in future wireless communication systems [35]. A potential solution is to adopt antenna selection schemes without using more resources in terms of hardware quality or extra antennas, to improve the system performance, which results in improvement of EE.
The contribution of each antenna to the signal and distortion power depends on the received power to that individual antenna. Antennas receiving more power may contribute more to the useful signal power at the cost of introducing more distortion. In Section III, we have seen how the receiver array size, given by the LIS radius, affects the performance of the whole system. We assumed that the area for signal reception was selected as a disk around the center of the array, assuming that the UE was also aligned with the CPL. The question that arises here would be: Can we improve the signal reception performance, i.e. SNDR, by selecting the antennas from another region? In other words, can we get an improvement from considering a reception area with a different shape than a disk with a center aligned with the UE position? This question can be formulated as an antenna selection problem, where we consider a LIS with $N\gg 1$ elements, and a resource constraint which forces the LIS to only use $N_{\text{max}}$ of the antenna elements for signal reception.
@@
$\max_{z_{1},\dots,z_{N}}$~ $\frac{P\sum_{n=1}^{N}z_{n}\left|\tilde{h}_{n}\right|^{2}}{\frac{\sum_{n=1}^{N}z_{n}C_{n}\left|\tilde{h}_{n}\right|^{2}}{\sum_{n=1}^{N}z_{n}\left|\tilde{h}_{n}\right|^{2}}+\sigma^{2}},$ (42)
@@
Let us define the set of selected antennas as $\mathcal{F}=\{i\in\mathbb{N}~|~r<x_{i}^{2}+y_{i}^{2}<R\}$. Therefore, we have $z_{i}=1$ if $i\in\mathcal{F}$. By leveraging Lemma III.1, the following approximation then holds for the numerator in (42),
@@
$\frac{\sum_{n=1}^{N}z_{n}C_{n}\left|\tilde{h}_{n}\right|^{2}}{\sum_{n=1}^{N}z_{n}\left|\tilde{h}_{n}\right|^{2}}=\frac{P\kappa\sum_{n\in\mathcal{F}}\left|{h}_{n}\right|^{4}}{\sum_{n\in\mathcal{F}}|h_{n}|^{2}}$ (44)
@@
which has the benefit of being a continuous function of only two variables, and it has a reduced complexity over (42) when solved by numerical methods. An algorithm to solve this problem numerically is introduced in Appendix B. Numerical results in Section VI show that solving the approximated optimization problem (45) results in negligible performance loss compared to directly solving (42).
@@ V LIS Panel Selection
$\max_{z_{n}}$~ $\frac{PM\sum_{n=1}^{N_{p}}z_{n}\left|\tilde{h}_{n}\right|^{2}}{\frac{\sum_{n=1}^{N_{p}}z_{n}C_{n}\left|\tilde{h}_{n}\right|^{2}}{\sum_{n=1}^{N_{p}}z_{n}\left|\tilde{h}_{n}\right|^{2}}+\sigma^{2}},$ (47)
@@
$2\beta|{a}_{3}|^{2}\rho^{4}+4\alpha|{a}_{3}|^{2}\rho^{3}-2\beta\sigma^{2}\rho-\alpha\sigma^{2}=0.$ (51)
@@
$\rho_{\text{opt}}=\sqrt[3]{-\frac{c_{0}}{2}+\sqrt{\Delta}}+\sqrt[3]{-\frac{c_{0}}{2}-\sqrt{\Delta}}.$ (55)